# The Spatial Complexity of Optical Computing and How to Reduce It

### v1

@@ I.1 The Spatial Complexity of Free-Space Optics
Figure 4: Space-efficient computing with free-space optics. Schematic of (a) conventional, (b) row sparse, and (c) local sparse ONNs performing a classification task on the fashion-MNIST dataset. (d-f) Thicknesses of the three interlayer regions in conventional (top row), row sparse (middle row), and local sparse (bottom row) ONNs. To demonstrate the trade-off between model thickness and accuracy, each model type is pruned using five different pruning thresholds (see Methods). Error bars represent one standard deviation of thickness (along $y$-axis) and accuracy (along $x$-axis) across eight different random seeds. A local sparse ONN with a pruning threshold of $\tau=0.01$, compared to a conventional ONN with $\tau=0.05$, drastically reduces the thickness while only compromising accuracy by $3.6\%$ on the fashion-MNIST dataset (vertical orange arrows; see Supplementary Note 6).
@@ I.3 Space-Efficient Computing on Photonic Chips: Analysis and Design
To evaluate the trade-off between $n_{\text{MZI}}$ and the model’s inference accuracy, we train networks with several degrees of block-diagonalization on the three MNIST-like datasets used in the previous section [31, 32, 33]. Each network has $5$ bias-free layers with $\left[784,100,100,10,10\right]$ neurons. Each layer, except the last one, is followed by a SiLU activation function [34]. We consider six different block-diagonalized architectures, named from “b-d1” to “b-d6”, where $n_{\text{MZI}}$ varies from approximately $3,200$ to approximately $22,600$. Fig. 5 summarizes the trade-off between $n_{\text{MZI}}$ and accuracy. These six models share one important feature, the weight (size $10\times 10$) connecting the last two layers remains unpruned, which we find is crucial for maintaining satisfactory accuracy. (Results on models with a pruned last weight are included in Supplementary Note 6.) Remarkably, model “b-d4”, with its largest block being only $10\times 10$ in size, achieves accuracies of $93.24\pm 0.40\%$, $84.51\pm 0.16\%$, and $66.76\pm 0.25\%$ on the MNIST, fashion-MNIST, Kuzushiji-MNIST datasets, respectively. Implementing “b-d4” using the MZI-grid architecture already requires fewer than $4,200$ MZIs, compared to the unpruned implementation’s approximately $0.3$ million MZIs—a reduction of about $98.7\%$. We also expect that emergent photonic components, for example, inverse-designed, programmable multiport couplers [39, 40, 41], could enable an even more compact and flexibly tunable on-chip implementation of similar models based on small, decoupled blocks.
Finally, we extend this block-diagonalization approach to 1. develop block-circulant ONNs, whose weight matrices consist of rows of cyclically shifting blocks, leveraging light diffraction for space-efficient computing (e.g.,[42]); and 2. prune models optimized for edge AI applications, e.g., MobileNetV2 [43] for CIFAR-10 [44] classification. We block-diagonalize the fully connected (FC) layer of MobileNetV2 and, with a largest block size of $10\times 10$, reduce $n_{\text{MZI}}$ by $\sim 99\%$ (considering the photonic implementation of FC layers only), while the accuracy decreases only slightly from $92.97\pm 0.24\%$ to $92.41\pm 0.25\%$. (See Supplementary Note 4 for both models’ configurations and performance.) With these extensions, we envision that a hybrid architecture—employing digital implementations for convolution layers, nonlinear activations, shortcut connections, etc., and photonic analog components for the classifier based on block-diagonal (or block-circulant) FC layers—could facilitate FLOP-, memory-, and space-efficient computation acceleration in real-world edge applications.
@@ II Discussion
III Data and Code Availability

### v2

@@ I.1 The Spatial Complexity of Free-Space Optics
Figure 4: Space-efficient computing with free-space optics. Schematic of (a) conventional, (b) row sparse, and (c) local sparse ONNs performing a classification task on the fashion-MNIST dataset. (d-f) Thicknesses of the three interlayer regions in conventional (top row), row sparse (middle row), and local sparse (bottom row) ONNs. To demonstrate the trade-off between model thickness and accuracy, each model type is pruned using five different pruning thresholds (see Methods). Error bars represent one standard deviation of thickness (along $y$-axis) and accuracy (along $x$-axis) across eight different random seeds. A local sparse ONN with a pruning threshold of $\tau=0.01$, compared to a conventional ONN with $\tau=0.05$, drastically reduces the thickness while only compromising accuracy by $3.6\%$ on the fashion-MNIST dataset (vertical orange arrows; see Supplementary Note 5).
@@ I.3 Space-Efficient Computing on Photonic Chips: Analysis and Design
To evaluate the trade-off between $n_{\text{MZI}}$ and the model’s inference accuracy, we train networks with several degrees of block-diagonalization on the three MNIST-like datasets used in the previous section [31, 32, 33]. Each network has $5$ bias-free layers with $\left[784,100,100,10,10\right]$ neurons. Each layer, except the last one, is followed by a SiLU activation function [34]. We consider six different block-diagonalized architectures, named from “b-d1” to “b-d6”, where $n_{\text{MZI}}$ varies from approximately $3,200$ to approximately $22,600$. Fig. 5 summarizes the trade-off between $n_{\text{MZI}}$ and accuracy. These six models share one important feature, the weight (size $10\times 10$) connecting the last two layers remains unpruned, which we find is crucial for maintaining satisfactory accuracy. (Results on models with a pruned last weight are included in Supplementary Note 5.) Remarkably, model “b-d4”, with its largest block being only $10\times 10$ in size, achieves accuracies of $93.24\pm 0.40\%$, $84.51\pm 0.16\%$, and $66.76\pm 0.25\%$ on the MNIST, fashion-MNIST, Kuzushiji-MNIST datasets, respectively. Implementing “b-d4” using the MZI-grid architecture already requires fewer than $4,200$ MZIs, compared to the unpruned implementation’s approximately $0.3$ million MZIs—a reduction of about $98.7\%$. We also expect that emergent photonic components, for example, inverse-designed, programmable multiport couplers [39, 40, 41], could enable an even more compact and flexibly tunable on-chip implementation of similar models based on small, decoupled blocks.
Finally, we extend this block-diagonalization approach to 1. develop block-circulant ONNs, whose weight matrices consist of rows of cyclically shifting blocks, leveraging light diffraction for space-efficient computing (e.g., [42]); and 2. prune models optimized for edge AI applications, e.g., MobileNetV2 [43] for CIFAR-10 [44] classification. We block-diagonalize the fully connected (FC) layer of MobileNetV2 and, with a largest block size of $10\times 10$, reduce $n_{\text{MZI}}$ by $\sim 99\%$ (considering the photonic implementation of FC layers only), while the accuracy decreases only slightly from $92.97\pm 0.24\%$ to $92.41\pm 0.25\%$. (See Supplementary Note 5 for the model architecture.) With these extensions, we envision that a hybrid architecture—employing digital implementations for convolution layers, nonlinear activations, shortcut connections, etc., and photonic analog components for the classifier based on block-diagonal (or block-circulant) FC layers—could facilitate FLOP-, memory-, and space-efficient computation acceleration in real-world edge applications.
@@ II Discussion
III Acknowledgements
This work was supported by the Air Force Office of Scientific Research with grant no. FA9550-22-1-0204 and Office of Naval Research with grant no. N00014-22-1-2486. Y. Li thanks Prof. Jon Kleinberg for helpful discussions on information networks.
IV Data and Code Availability